# Galactic Orbits of Interstellar Objects

### v1

@@ Abstract
Context. The first interstellar objects, such as ‘Oumuamua, Borisov and IM1, were discovered over the past decade.
@@
Results. The small vertical extent of ‘Oumuamua’s past trajectory suggests that it originated near the mid-plane of the thin disk, implying a likely age younger than 1–2 Gyr. The maximal excursion of the comet Borisov is similar to that of the Sun, suggesting a similar age. The meteor IM1 exhibits yet larger vertical excursions, suggesting an older source. Finally, we show that human-made interstellar probes, like Voyager 1 or Pioneer 10 will arrive at the opposite side of the Milky Way disk relative to the Sun in $\sim 2$Gyr and return to the vicinity of the Sun before it becomes a red giant.
@@ 1 Introduction
The discovery of interstellar objects like ‘Oumuamua and Borisov over the past decade has sparked significant interest in understanding their origins and dynamics (see reviews by Siraj & Loeb (2022a), Jewitt and Seligman (2023), Jewitt (2024), and references therein). A fundamental unknown is the likely source of each of these objects (Bailer-Jones et al., 2018, 2020). Constraining the sources could shed light on the nature of these interstellar objects and the astrophysical processes that created them (see, for example, Zhang & Lin (2020) or Loeb & MacLeod (2024)).
@@
We also apply the same numerical approach to calculating the future trajectories of the interstellar probes launched by NASA decades ago, Voyager 1 & 2 and Pioneer 10 & 11.
The organization of this paper is as follows. Section 2 describes our method of calculation and the resulting Galactic trajectories for the individual interstellar objects, with subsections dedicated to Borisov (§2.2), ‘Oumuamua (§2.3) and IM1 (§2.4). In Section 3, we use the resulting vertical excursion of these trajectories to derive the probability distribution for the likely age for each of these objects. Section 5 considers the future trajectories of human-made interstellar probes. Finally, we summarize the implications of our results in Section 6.
@@ 2.1 Method of Calculation
In order to convert the velocity measurements in the Solar system to the Galactic frame of reference, we incorporate the motion of the Sun relative to the Local Standard of Rest (LSR), with the Galactic components: $U_{\odot}=10.79\pm 0.56$ km s-1, $V_{\odot}=11.06\pm 0.94$ km s-1, and $W_{\odot}=7.66\pm 0.43$ km s-1 (Robin et al., 2022). Before entering the solar system, Borisov’s velocity was $(U,V,W)=(33.1,-6.8,8.3)$ km s-1, while ‘Oumuamua’s velocity relative to the Sun is $(U-U_{\odot},V-V_{\odot},W-W_{\odot})=(-11.457\pm 0.009,-22.395\pm 0.009,-7.74\pm 0.011)$ km s-1 (Bailer-Jones et al., 2020; Mamajek, 2017). For the interstellar meteor, IM1, we adopt $(U,V,W)=(32.7\pm 5.8,-4.5\pm 1.5,26.1\pm 2.0)$ km s-1 (Siraj & Loeb, 2022b).
2.2 Comet Borisov
The past evolution of the distance of the interstellar comet Borisov from the Sun is shown in figure 1. The Sun-Borisov separation follows a period of about 3.2 Gyr, with Borisov being on the other side of the Milky-Way disk relative to the Sun about 1.6 Gyr ago.
Figure 1: Distance of the comet Borisov from the Sun back in time with the present time represented by 0.
The radial ($R$) and vertical ($z$) extent of Borisov’s trajectory relative to the Galactic plane over the past 10 Gyr are depicted in figure 2 (orange) and resemble the corresponding ranges for the Sun (blue).
Figure 2: Trajectory of the comet Borisov in the $R-z$ and $x-y$ plane of the Milky-Way disk.
2.3 ’Oumuamua
@@ 2.2 Comet Borisov
The past evolution of the distance of the interstellar object ‘Oumuamua from the Sun is shown in figure 3.The ‘Oumuamua-Sun separation follows a period of about 2.2 Gyr,with ‘Oumuamua being on the other side of the Milky Way disk relative to the Sun about 1.1 Gyr ago.
Figure 3: Distance of ’Oumuamua from the Sun back in time.
@@
The radial ($R$) and vertical ($z$) extent of ‘Oumuamua’s trajectory relative to the Galactic plane over the past 10 Gyr are depicted in figure 4 (orange). As a result of its low vertical velocity relative to the LSR, the vertical excursion of ‘Oumuamua is a factor of $\sim 6$ smaller than that of the Sun (blue).
@@
Figure 4: Trajectory of ’Oumuamua in the $R-z$ and the $x-y$ plane of the Mily-Way disk.
2.4 IM1
@@
The past evolution of the distance of the interstellar meteor IM1 from the Sun is shown in figure 5. The IM1-Sun separation follows a period of about 4.3 Gyr,with IM1 being on the other side of the Milky Way disk relative to the Sun about 2.15 Gyr ago.
@@
Figure 5: Distance of IM1 from the Sun back in time.
@@
The radial ($R$) and vertical ($z$) extent of IM1’s trajectory relative to the Galactic plane over the past 10 Gyr are depicted in figure 6 (orange). As a result of its high vertical velocity, the vertical excursion of IM1 is $\sim 2.5$ times larger that of the Sun (blue).
Figure 6: Trajectory of IM1 in the $R-z$ and the Galactic $x-y$ plane.
@@ 3.2 Likelihood of Ages for individual Interstellar Objects
Figure 8 shows the probability distribution of likely ages for the comet Borisov, which straddles the age of the Sun at $4.6$ Gyr.
Figure 8: Probability distribution of age, $p(t)$ in units of Gyr-1, for the interstellar comet Borisov with $z_{\text{max}}=0.12$ kpc.
In contrast, ‘Oumuamua is much younger with a likely age of 1-2 Gyr, based on its much smaller value of $z_{\rm max}$, as shown in Figure 9.
Figure 9: Probability distribution of age, $p(t)$ in units of Gyr-1, for Oumuamua with $z_{\text{max}}=0.015$ kpc.
Finally, IM1 follows the probability distribution of the entire Galactic disk population, SFR(t), because of its large value of $z_{\rm max}$, as shown in Figure 10.
Figure 10: Probability distribution of age, $p(t)$ in units of Gyr-1, for IM1 with $z_{\text{max}}=0.45$ kpc.
4 Interstellar Probes
Finally, we derive the future trajectories of the first four interstellar probes, Voyager 1 & 2, Pioneer 10 & 11, which are making their way out of the Solar system. For each probe, we adopt the current velocity relative to the Sun based on the following directions of motion (right ascension and declination) and speeds (Bailer-Jones & Farnocchia, 2019):
  • Voyager 1: $\alpha=262.8760^{\circ}$, $\delta=12.3199^{\circ}$, $v=16.6048\,\text{km s}^{-1}$.
  • Voyager 2: $\alpha=316.2717^{\circ}$, $\delta=-67.5491^{\circ}$, $v=14.8550\,\text{km s}^{-1}$.
  • Pioneer 10: $\alpha=83.4169^{\circ}$, $\delta=26.2171^{\circ}$, $v=11.3149\,\text{km s}^{-1}$.
  • Pioneer 11: $\alpha=291.8277^{\circ}$, $\delta=-9.2212^{\circ}$, $v=10.4439\,\text{km s}^{-1}$.
We transformed the velocity vectors to Cartesian Galactic coordinates and added the velocity vector of the Sun. Below we describe the resulting trajectory for each of the probes.
4.1 Voyager 1
The future evolution of the distance of the Voyager 1 from the Sun is shown in figure 11. The separation follows a period of about 3.6 Gyr, with Voyager reaching the other side of the Milky Way disk relative to the Sun in about 1.8 Gyr.
Figure 11: Distance of Voyager 1 from the Sun as a function of future time, with 0 representing the present time.
The radial ($R$) and vertical ($z$) extent of Voyager 1’s trajectory relative to the Galactic plane in the next 10 Gyr are depicted in figure 12 (orange) and resemble the corresponding ranges for the Sun (blue).
Figure 12: Trajectories of Voyager 1 in the $R-z$ and the Galactic $x-y$ plane.
4.2 Voyager 2
The future evolution of the distance of the Voyager 2 from the Sun is shown in figure 13. The Voyager 2-Sun separation follows a period of about 13 Gyr, comparable to the age of the Universe, with the probe reaching the other side of the Milky Way disk relative to the Sun in about 6.5 Gyr. Interestingly, Voyager 2 is predicted to arrive within 0.1kpc from the Sun on its first return.
Figure 13: Distance evolution of Voyager 2 from the Sun.
The radial ($R$) and vertical ($z$) extent of Voyager 2’s trajectory relative to the Galactic plane in the next 10 Gyr are depicted in figure 14 (orange) and resemble the corresponding ranges for the Sun (blue).
Figure 14: Trajectories of Voyager 2 in the $R-z$ and the Galactic $x-y$ plane.
4.3 Pioneer 10
The future evolution of the distance of the Pioneer 10 from the Sun is shown in figure 15. The separation follows a period of about 4.6 Gyr, with Pioneer 10 reaching the other side of the Milky Way disk relative to the Sun in about 2.3 Gyr.
Figure 15: Distance evolution of Pioneer 10 from the Sun.
The radial ($R$) and vertical ($z$) extent of Pioneer 10’s trajectory relative to the Galactic plane in the next 10 Gyr are depicted in figure 16 (orange) and resemble the corresponding ranges for the Sun (blue).
Figure 16: Trajectories of Pioneer 10 in the $R-z$ and the Galactic $x-y$ plane.
4.4 Pioneer 11
The future evolution of the distance of the Pioneer 10 from the Sun is shown in figure 17. The separation follows a period of about 5 Gyr, with Pioneer 12 reaching the other side of the Milky Way disk relative to the Sun in about 2.5 Gyr.
Figure 17: Distance evolution of Pioneer 11 from the Sun.
The radial ($R$) and vertical ($z$) extent of Pioneer 10’s trajectory relative to the Galactic plane in the next 10 Gyr are depicted in figure 18 (orange) and resemble the corresponding ranges for the Sun (blue).
Figure 18: Trajectories of Pioneer 11 in the $R-z$ and the Galactic $x-y$ plane.
@@ 5 Conclusions
We have integrated the Galactic trajectories of known interstellar objects back in time. The vertical extent of the trajectory of ’Oumuamua suggests that it originated near the disk mid-plane of the thin disk, where the ages of most stars are below 1–2 Gyr. In contrast, the comet Borisov’s vertical excursion is similar to that of the Sun, suggesting a similar age. Given its highest vertical excursions, the putative source star of IM1 is most likely the oldest.
@@
In addition, our calculations indicate that the interstellar probes Voyager 1 and Pioneer 10 will reach the opposite side of the Milky-Way disk relative to the Sun in $\sim 2$ Gyr from now, and will return to the Solar vicinity in $\sim 4$ Gyr. This future return will occur long before the Sun will become a red giant star in $\sim 7.6$ Gyr (Schröder & Smith, 2008).
Acknowledgements. We thank Morgan MacLeod for helpful comments on the manuscript. This work was supported in part by Harvard’s Institute for Theory & Computation, the Harvard Astronomy and Physics departments,the BHI and the Galileo Project.

### v4

# Galactic Trajectories of Interstellar Objects 1I/’Oumuamua, 2I/Borisov, and 3I/Atlas

@@ Abstract
Context. The first interstellar objects, 1I/‘Oumuamua, 2I/Borisov and 3I/ATLAS, were discovered over the past decade.
@@
Results. The small vertical extent of 1I/‘Oumuamua’s past trajectory suggests that it originated near the mid-plane of the thin Galactic disk of stars, implying a likely age younger than 1–2 Gyr. The maximal excursion of the comet 2I/Borisov is similar to that of the Sun, suggesting a similar age. The interstellar object 3I/ATLAS exhibits significantly larger vertical excursions, with $z_{\text{max}}=0.475$ , suggesting that it originated from an older source population in the thick Galactic disk.
@@ 1 Introduction
The discovery of interstellar objects over the past decade has sparked significant interest in understanding their origins and dynamics (see reviews by Siraj & Loeb (2022a), Jewitt and Seligman (2023), Jewitt (2024), and references therein). A fundamental unknown is the likely origin of each of these objects (Bailer-Jones et al., 2018, 2020). Constraining the sources could shed light on the nature of these interstellar objects and the astrophysical processes that created them (see, for example, Zhang & Lin (2020) or Loeb & MacLeod (2024)).
@@
The organization of this paper is as follows. Section 2 describes our method of calculation and the resulting Galactic trajectories for the individual interstellar objects, with subsections dedicated to 2I/Borisov (§2.2), 1I/‘Oumuamua (§2.3) and 3I/ATLAS (§2.4). In Section 3, we use the resulting vertical excursion of these trajectories to derive the probability distribution for the likely age for each of these objects. Finally, we summarize the implications of our results in Section 4.
@@ 2.1 Method of Calculation
In order to convert the velocity measurements in the Solar system to the Galactic frame of reference, we incorporate the motion of the Sun relative to the Local Standard of Rest (LSR), with the Galactic components: $U_{\odot}=10.79\pm 0.56$ km s-1, $V_{\odot}=11.06\pm 0.94$ km s-1, and $W_{\odot}=7.66\pm 0.43$ km s-1 (Robin et al., 2022). Before entering the solar system, 2I/Borisov’s velocity was $(U,V,W)=(33.1,-6.8,8.3)$ km s-1, while 1I/‘Oumuamua’s velocity relative to the Sun is $(U-U_{\odot},V-V_{\odot},W-W_{\odot})=(-11.457\pm 0.009,-22.395\pm 0.009,-7.74\pm 0.011)$ km s-1 (Bailer-Jones et al., 2020; Mamajek, 2017). For the interstellar object 3I/ATLAS (C/2025 N1), we adopt its inbound heliocentric velocities of $(U,V,W)=(-54.4,-20.3,+19.5)$ km s-1 and corresponding Galactocentric cylindrical velocities of $(v_{R},v_{z},v_{\phi})=(+43.62,+27.18,+223.76)$ km s-1, derived from JPL orbital solution JPL#5. We verified the orbital calculation accuracy through direct API comparison with JPL Horizons (achieving $<10^{-10}$ AU precision), independent mathematical validation of the elements-to-Cartesian conversion, and consistency checks with predicted planetary encounter geometries. The Galactocentric velocities were obtained via standard coordinate transformations incorporating solar motion parameters.
2.2 1I/’Oumuamua
@@
The past evolution of the distance of the interstellar object 1I/‘Oumuamua from the Sun is shown in figure 1. The 1I/‘Oumuamua-Sun separation follows a period of about 2.2 Gyr, with 1I/‘Oumuamua being on the other side of the Milky Way disk relative to the Sun about 1.1 Gyr ago.
Figure 1: Distance of 1I/’Oumuamua from the Sun back in time.
@@
The radial ($R$) and vertical ($z$) extent of 1I/‘Oumuamua’s trajectory relative to the Galactic plane are depicted in figure 2 for both 5 Gyr (left panel) and 1 Gyr (right panel) integrations. As a result of its low vertical velocity relative to the LSR, the vertical excursion of 1I/‘Oumuamua (orange) is a factor of $\sim 6$ smaller than that of the Sun (blue).
@@
Figure 2: Trajectory of 1I/’Oumuamua in the $R-z$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
The motion in the Galactic plane ($x-y$) is shown in figure 3, comparing 5 Gyr and 1 Gyr integrations.
Figure 3: Trajectory of 1I/’Oumuamua in the $x-y$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
2.3 2I/Borisov
@@
The past evolution of the distance of the interstellar comet 2I/Borisov from the Sun is shown in figure 4. The Sun-2I/Borisov separation follows a period of about 3.2 Gyr, with 2I/Borisov being on the other side of the Milky-Way disk relative to the Sun about 1.6 Gyr ago.
@@
Figure 4: Distance of the comet 2I/Borisov from the Sun back in time with the present time represented by 0.
The radial ($R$) and vertical ($z$) extent of 2I/Borisov’s trajectory relative to the Galactic plane are depicted in figure 5 for both 5 Gyr (left panel) and 1 Gyr (right panel) integrations. The trajectories of 2I/Borisov (orange) resemble the corresponding ranges for the Sun (blue).
Figure 5: Trajectory of the comet 2I/Borisov in the $R-z$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
The motion in the Galactic plane ($x-y$) is shown in figure 6, again comparing 5 Gyr and 1 Gyr integrations.
Figure 6: Trajectory of the comet 2I/Borisov in the $x-y$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
2.4 3I/ATLAS
The past evolution of the distance of the interstellar object 3I/ATLAS from the Sun is shown in figure 7. The 3I/ATLAS-Sun separation follows a period of about 3.63 Gyr, with 3I/ATLAS being on the other side of the Milky Way disk relative to the Sun about 1.815 Gyr ago.
Figure 7: Distance of 3I/ATLAS from the Sun back in time.
@@
The radial ($R$) and vertical ($z$) extent of 3I/ATLAS’s trajectory relative to the Galactic plane are depicted in figure 8 for both 5 Gyr (left panel) and 1 Gyr (right panel) integrations. As a result of its vertical velocity, the vertical excursion of 3I/ATLAS (orange) is larger than that of the Sun (blue).
Figure 8: Trajectory of 3I/ATLAS in the $R-z$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
The motion in the Galactic plane ($x-y$) is shown in figure 9, comparing 5 Gyr and 1 Gyr integrations.
Figure 9: Trajectory of 3I/ATLAS in the $x-y$ plane of the Milky-Way disk. Left panel shows 5 Gyr integration, right panel shows 1 Gyr integration.
@@ 3.2 Likelihood of Ages for individual Interstellar Objects
Figure 11 shows the probability distribution of likely ages for the comet 2I/Borisov, which straddles the age of the Sun at $4.6$ Gyr.
Figure 11: Probability distribution of age, $p(t)$ in units of Gyr-1, for the interstellar comet 2I/Borisov with $z_{\text{max}}=0.12$ kpc.
In contrast, 1I/‘Oumuamua is much younger with a likely age of 1-2 Gyr, based on its much smaller value of $z_{\rm max}$, as shown in Figure 12.
Figure 12: Probability distribution of age, $p(t)$ in units of Gyr-1, for 1I/‘Oumuamua with $z_{\text{max}}=0.015$ kpc.
Finally, 3I/ATLAS may follow a probability distribution similar to the entire Galactic disk population because of its potentially large value of $z_{\rm max}$, as shown in Figure 13.
Figure 13: Probability distribution of age, $p(t)$ in units of Gyr-1, for 3I/ATLAS with $z_{\text{max}}=\text{0.475}$ kpc.
@@ 4 Conclusions
We have integrated the Galactic trajectories of known interstellar objects back in time. The vertical extent of the trajectory of 1I/’Oumuamua suggests that it originated near the disk mid-plane of the thin disk, where the ages of most stars are below 1–2 Gyr. In contrast, the comet 2I/Borisov’s vertical excursion is similar to that of the Sun, suggesting a similar age. Given its high vertical excursions, the putative source star of 3I/ATLAS is likely older than that of 1I/’Oumuamua or 2I/Borisov.
@@
Concluding note: As we were about to submit our manuscript for publication, Hopkins et al. (2025) posted a paper which addresses some of the same topics. Our main results are consistent with theirs.
Acknowledgements. We thank Matthew Hopkins for his careful cross-checking of our paper. This work was supported in part by Harvard University, the Black Hole Initiative and the Galileo Project.